(Common stock valuation) The common stock of NCP paid $1.25 in dividends last year. Dividends are expected to grow at an annual rate of 9.80 percent for an indefinite number of years.

a. If NCP’s current market price is $20.35 per share, what is the stock’s expected rate of return?
b. If your required rate of return is 11.8 percent, what is the value of the stock for you?
c. Should you make the investment?
Answer #1

a. Expected rate of return = (D1 / P0) + g = [$1.25(1.098) / $20.35] + 0.098 = 16.54%

b. Value of the stock = D1 / (r - g) = $1.25(1.098) / (0.118 - 0.098) = $68.63

c. Yes